OpenStudy (anonymous):

49-x^2/7x-1 divided by 7-x/5x^7

OpenStudy (anonymous):

\[(49-x ^{2}\div7x-1)\div(7-x \div5x ^{7})\]

OpenStudy (anonymous):

Do you have a set of answers?

OpenStudy (anonymous):

nope! but i know you do the reciprocal of the denominator fraction so it cancels out. so then you're left with \[49-x^2\div7x-1\]

OpenStudy (anonymous):

This is what I have. Though I'm not too sure it is correct. \[\frac{ 49-x ^{2} }{ 7x-1 }\times \frac{ 5x ^{7} }{ 7-x }\]\[\frac{ -(x+7)(x-7)(5x ^{7}) }{ (7x-1)(7-x) }\] the (x-7) cancels out leaving.\[\frac{ -(x+7)(5x ^{7}) }{ (7x-1) }\]\[\frac{5x ^{7}(-x-7) }{ (7x-1) }\] I can't get any further than that. I'm not sure if this is correct.

OpenStudy (anonymous):

Wait. I found an error.

OpenStudy (anonymous):

Sorry I couldn't explain the process, but I looked online and the answer is x=-7. http://www.wolframalpha.com/input/?i=Solve+%28%2849-x%5E2%29%2F%287x-1%29%29%2F%28%287-x%29%2F%285x%5E7%29%29